# Convergent Privacy Framework with Contractive GNN Layers for Multi-hop Aggregations

### v1

@@ E.2. Accuracy of $\mathsf{CARIBOU}$-NDP
Dataset Chain-S Chain-M Chain-L Chain-X Computers Facebook PubMed Cora Photo $\epsilon=1$ $\mathsf{CARIBOU}$ $78.1\%$ $80.0\%$ $70.0\%\mathrel{\hskip 5.12128pt}$ $65.0\%\mathrel{\hskip 5.12128pt}$ $92.0\%$ $74.0\%\mathrel{\hskip 5.12128pt}$ $87.9\%$ $84.3\%$ $95.6\%$ DPDGC $43.8\%$blacktriangledown $50.0\%$blacktriangledown $51.7\%$blacktriangledown $39.0\%$blacktriangledown $88.0\%$ $60.6\%\mathrel{\hskip 5.12128pt}$ $88.3\%$ $75.5\%$ $92.5\%$ GAP $65.6\%\mathrel{\hskip 5.12128pt}$ $67.5\%\mathrel{\hskip 5.12128pt}$ $61.7\%\mathrel{\hskip 5.12128pt}$ $55.5\%\mathrel{\hskip 5.12128pt}$ $87.0\%$ $68.3\%\mathrel{\hskip 5.12128pt}$ $87.2\%$ $76.0\%$ $92.8\%$ SAGE $53.1\%$blacktriangledown $45.0\%$blacktriangledown $53.3\%$blacktriangledown $51.0\%$blacktriangledown $77.8\%$ $48.2\%$blacktriangledown $85.0\%$ $60.0\%$ $82.4\%$ $\epsilon=2$ $\mathsf{CARIBOU}$ $90.6\%$ $82.5\%$ $71.7\%$ $68.0\%$ $92.0\%$ $73.8\%\mathrel{\hskip 5.12128pt}$ $89.1\%$ $86.0\%$ $95.7\%$ DPDGC $43.8\%$blacktriangledown $50.0\%$blacktriangledown $51.7\%$blacktriangledown $39.0\%$blacktriangledown $88.2\%$ $66.7\%\mathrel{\hskip 5.12128pt}$ $88.2\%$ $77.5\%$ $93.3\%$ GAP $65.6\%\mathrel{\hskip 5.12128pt}$ $67.5\%\mathrel{\hskip 5.12128pt}$ $61.7\%\mathrel{\hskip 5.12128pt}$ $55.0\%$blacktriangledown $88.0\%$ $71.7\%\mathrel{\hskip 5.12128pt}$ $87.3\%$ $76.8\%$ $93.3\%$ SAGE $43.8\%$blacktriangledown $45.0\%$blacktriangledown $53.3\%$blacktriangledown $51.0\%$blacktriangledown $76.1\%$ $48.1\%$blacktriangledown $84.6\%$ $60.1\%$ $82.4\%$ $\epsilon=4$ $\mathsf{CARIBOU}$ $90.6\%$ $82.5\%$ $71.7\%$ $65.0\%\mathrel{\hskip 5.12128pt}$ $92.2\%$ $73.9\%\mathrel{\hskip 5.12128pt}$ $89.5\%$ $86.7\%$ $95.8\%$ DPDGC $43.8\%$blacktriangledown $50.0\%$blacktriangledown $51.7\%$blacktriangledown $39.0\%$blacktriangledown $88.9\%$ $73.3\%\mathrel{\hskip 5.12128pt}$ $88.4\%$ $75.1\%$ $94.2\%$ GAP $62.5\%\mathrel{\hskip 5.12128pt}$ $60.0\%\mathrel{\hskip 5.12128pt}$ $61.7\%\mathrel{\hskip 5.12128pt}$ $55.5\%\mathrel{\hskip 5.12128pt}$ $88.8\%$ $73.6\%\mathrel{\hskip 5.12128pt}$ $87.7\%$ $76.9\%$ $93.8\%$ SAGE $50.0\%$blacktriangledown $47.5\%$blacktriangledown $51.7\%$blacktriangledown $54.0\%$blacktriangledown $79.1\%$ $50.3\%$blacktriangledown $85.8\%$ $63.3\%$ $85.7\%$ $\epsilon=8$ $\mathsf{CARIBOU}$ $81.2\%$ $82.5\%$ $73.3\%$ $68.0\%\mathrel{\hskip 5.12128pt}$ $92.2\%$ $74.2\%\mathrel{\hskip 5.12128pt}$ $89.7\%$ $87.6\%$ $95.8\%$ DPDGC $43.8\%$blacktriangledown $50.0\%$blacktriangledown $51.7\%$blacktriangledown $39.0\%$blacktriangledown $89.4\%$ $78.6\%\mathrel{\hskip 5.12128pt}$ $88.6\%$ $76.0\%$ $94.6\%$ GAP $59.4\%\mathrel{\hskip 5.12128pt}$ $60.0\%\mathrel{\hskip 5.12128pt}$ $61.7\%\mathrel{\hskip 5.12128pt}$ $55.5\%\mathrel{\hskip 5.12128pt}$ $89.6\%$ $75.0\%\mathrel{\hskip 5.12128pt}$ $88.0\%$ $78.2\%$ $94.6\%$ SAGE $43.8\%$blacktriangledown $47.5\%$blacktriangledown $51.7\%$blacktriangledown $61.0\%\mathrel{\hskip 5.12128pt}$ $87.9\%$ $75.6\%\mathrel{\hskip 5.12128pt}$ $84.8\%$ $75.7\%$ $92.2\%$ $\epsilon=16$ $\mathsf{CARIBOU}$ $87.5\%$ $85.0\%$ $70.0\%\mathrel{\hskip 5.12128pt}$ $68.0\%$ $92.1\%$ $74.3\%\mathrel{\hskip 5.12128pt}$ $89.8\%$ $88.0\%$ $95.9\%$ DPDGC $43.8\%$blacktriangledown $50.0\%$blacktriangledown $51.7\%$blacktriangledown $40.0\%$blacktriangledown $90.4\%$ $81.2\%\mathrel{\hskip 5.12128pt}$ $88.9\%$ $77.7\%$ $93.8\%$ GAP $59.4\%\mathrel{\hskip 5.12128pt}$ $70.0\%\mathrel{\hskip 5.12128pt}$ $61.7\%\mathrel{\hskip 5.12128pt}$ $54.3\%$blacktriangledown $90.0\%$ $76.0\%\mathrel{\hskip 5.12128pt}$ $88.5\%$ $80.3\%$ $94.6\%$ SAGE $43.8\%$blacktriangledown $47.5\%$blacktriangledown $51.7\%$blacktriangledown $51.0\%$blacktriangledown $90.9\%$ $79.5\%\mathrel{\hskip 5.12128pt}$ $87.6\%$ $84.7\%$ $94.1\%$ $\epsilon=32$ $\mathsf{CARIBOU}$ $90.6\%$ $82.5\%$ $78.3\%$ $67.0\%$ $92.2\%$ $73.9\%\mathrel{\hskip 5.12128pt}$ $89.8\%$ $88.2\%$ $95.8\%$ DPDGC $43.8\%$blacktriangledown $50.0\%$blacktriangledown $51.7\%$blacktriangledown $40.0\%$blacktriangledown $91.3\%$ $82.9\%\mathrel{\hskip 5.12128pt}$ $88.8\%$ $79.9\%$ $94.3\%$ GAP $59.4\%\mathrel{\hskip 5.12128pt}$ $65.0\%\mathrel{\hskip 5.12128pt}$ $61.7\%\mathrel{\hskip 5.12128pt}$ $54.3\%$blacktriangledown $90.2\%$ $76.5\%\mathrel{\hskip 5.12128pt}$ $88.7\%$ $82.2\%$ $94.6\%$ SAGE $43.8\%$blacktriangledown $47.5\%$blacktriangledown $51.7\%$blacktriangledown $51.0\%$blacktriangledown $90.6\%$ $79.9\%\mathrel{\hskip 5.12128pt}$ $86.9\%$ $85.2\%$ $94.4\%$ Plain $\mathsf{CARIBOU}$ $100.0\%\mathrel{\hskip 5.12128pt}$ $100.0\%\mathrel{\hskip 5.12128pt}$ $100.0\%\mathrel{\hskip 5.12128pt}$ $100.0\%\mathrel{\hskip 5.12128pt}$ $92.2\%$ $78.1\%\mathrel{\hskip 5.12128pt}$ $89.9\%$ $88.4\%$ $95.8\%$ DPDGC $59.4\%\mathrel{\hskip 5.12128pt}$ $77.5\%\mathrel{\hskip 5.12128pt}$ $63.3\%\mathrel{\hskip 5.12128pt}$ $73.0\%\mathrel{\hskip 5.12128pt}$ $92.8\%$ $86.4\%\mathrel{\hskip 5.12128pt}$ $88.1\%$ $83.9\%$ $96.2\%$ GAP $97.9\%\mathrel{\hskip 5.12128pt}$ $100.0\%\mathrel{\hskip 5.12128pt}$ $93.3\%\mathrel{\hskip 5.12128pt}$ $100.0\%\mathrel{\hskip 5.12128pt}$ $90.9\%$ $78.3\%\mathrel{\hskip 5.12128pt}$ $89.1\%$ $85.1\%$ $95.2\%$ SAGE $59.4\%\mathrel{\hskip 5.12128pt}$ $55.0\%$blacktriangledown $60.0\%\mathrel{\hskip 5.12128pt}$ $58.0\%\mathrel{\hskip 5.12128pt}$ $91.6\%$ $79.7\%\mathrel{\hskip 5.12128pt}$ $87.0\%$ $82.1\%$ $94.2\%$ MLP $40.6\%$blacktriangledown $47.5\%$blacktriangledown $46.7\%$blacktriangledown $51.0\%$blacktriangledown $83.9\%$ $51.1\%$blacktriangledown $84.9\%$ $72.9\%$ $91.1\%$
Table 10. Top Accuracy of Node-level Privacy.
Dataset Chain-S Chain-M Chain-L Chain-X Cora Photo
$\epsilon=1$ $75.0\%$ $70.0\%$ $60.0\%$ $60.0\%$ $66.8\%$ $93.7\%$
$\epsilon=2$ $78.1\%$ $70.0\%$ $61.7\%$ $61.0\%$ $70.3\%$ $94.7\%$
$\epsilon=4$ $78.1\%$ $70.0\%$ $63.3\%$ $61.0\%$ $75.5\%$ $95.0\%$
$\epsilon=8$ $78.1\%$ $72.5\%$ $63.3\%$ $62.0\%$ $79.9\%$ $95.3\%$
$\epsilon=16$ $75.0\%$ $72.5\%$ $68.3\%$ $63.0\%$ $82.1\%$ $95.3\%$
$\epsilon=32$ $71.9\%$ $72.5\%$ $73.3\%$ $64.0\%$ $84.7\%$ $95.5\%$
E.3. More Results of EDP
Table 9 shows the mean accuracy compared between $\mathsf{CARIBOU}$ and other baselines for all datasets.
Figure 7. Accuracy of non-private $\mathsf{CARIBOU}$ and GAP under different $K$.
Figure 8. Ablation study of $K$ on $\mathsf{CARIBOU}$ (colored boxes) and GAP (blue lines). For each pair of $\epsilon$ and dataset (Chain-S, Chain-M and Chain-L), different $K$ are used.
Figure 9. Ablation study of $\epsilon$ on $\mathsf{CARIBOU}$ (colored boxes) and GAP (blue lines). For each pair of $K$ and dataset (Chain-S, Chain-M and Chain-L), different $\epsilon$ are used.
E.4. Choices of Hyper-parameters
Appendix F More Related Works
Long-range graphs and GNNs.
Despite GNNs’ success in structured data, modeling long-range dependencies remains challenging due to over-smoothing (Chen et al., 2020a), vanishing gradients, and exponentially growing receptive fields. Enforcing DP in long-range GNNs is particularly challenging, as these models aggregate node embeddings over multiple hops and broader neighborhoods. Accordingly, injecting DP noise must account for multi-hop aggregation, requiring careful design to balance privacy and utility.
Differential privacy for graph structures.

### v2

@@ A-F3 More Results
Dataset Computers Facebook PubMed Cora Photo Chain-S Chain-M Chain-L Chain-X EDP $\epsilon=1$ $\mathsf{CARIBOU}$ $92.0\%$ $74.0\%$ $87.9\%$ $84.3\%$ $95.6\%$ $78.1\%$ $80.0\%$ $70.0\%$ $65.0\%$ DPDGC $88.0\%$ $60.6\%$ $88.3\%$ $75.5\%$ $92.5\%$ $43.8\%$ $50.0\%$ $51.7\%$ $39.0\%$ GAP $87.0\%$ $68.3\%$ $87.2\%$ $76.0\%$ $92.8\%$ $65.6\%$ $67.5\%$ $61.7\%$ $55.5\%$ PertGraph $77.8\%$ $48.2\%$ $85.0\%$ $60.0\%$ $82.4\%$ $53.1\%$ $45.0\%$ $53.3\%$ $51.0\%$ $\epsilon=2$ $\mathsf{CARIBOU}$ $92.0\%$ $73.8\%$ $89.1\%$ $86.0\%$ $95.7\%$ $90.6\%$ $82.5\%$ $71.7\%$ $68.0\%$ DPDGC $88.2\%$ $66.7\%$ $88.2\%$ $77.5\%$ $93.3\%$ $43.8\%$ $50.0\%$ $51.7\%$ $39.0\%$ GAP $88.0\%$ $71.7\%$ $87.3\%$ $76.8\%$ $93.3\%$ $65.6\%$ $67.5\%$ $61.7\%$ $55.0\%$ PertGraph $76.1\%$ $48.1\%$ $84.6\%$ $60.1\%$ $82.4\%$ $43.8\%$ $45.0\%$ $53.3\%$ $51.0\%$ $\epsilon=4$ $\mathsf{CARIBOU}$ $92.2\%$ $73.9\%$ $89.5\%$ $86.7\%$ $95.8\%$ $90.6\%$ $82.5\%$ $71.7\%$ $65.0\%$ DPDGC $88.9\%$ $73.3\%$ $88.4\%$ $75.1\%$ $94.2\%$ $43.8\%$ $50.0\%$ $51.7\%$ $39.0\%$ GAP $88.8\%$ $73.6\%$ $87.7\%$ $76.9\%$ $93.8\%$ $62.5\%$ $60.0\%$ $61.7\%$ $55.5\%$ PertGraph $79.1\%$ $50.3\%$ $85.8\%$ $63.3\%$ $85.7\%$ $50.0\%$ $47.5\%$ $51.7\%$ $54.0\%$ $\epsilon=8$ $\mathsf{CARIBOU}$ $92.2\%$ $74.2\%$ $89.7\%$ $87.6\%$ $95.8\%$ $81.2\%$ $82.5\%$ $73.3\%$ $68.0\%$ DPDGC $89.4\%$ $78.6\%$ $88.6\%$ $76.0\%$ $94.6\%$ $43.8\%$ $50.0\%$ $51.7\%$ $39.0\%$ GAP $89.6\%$ $75.0\%$ $88.0\%$ $78.2\%$ $94.6\%$ $59.4\%$ $60.0\%$ $61.7\%$ $55.5\%$ PertGraph $87.9\%$ $75.6\%$ $84.8\%$ $75.7\%$ $92.2\%$ $43.8\%$ $47.5\%$ $51.7\%$ $61.0\%$ $\epsilon=16$ $\mathsf{CARIBOU}$ $92.1\%$ $74.3\%$ $9.8\% 88$$8.0\%$ $95.9\%$ $87.5\%$ $85.0\%$ $70.0\%$ $68.0\%$ DPDGC $90.4\%$ $1.2\% 88$$8.9\%$ $77.7\%$ $93.8\%$ $43.8\%$ $50.0\%$ $51.7\%$ $40.0\%$ GAP $90.0\%$ $76.0\%$ $88.5\%$ $80.3\%$ $94.6\%$ $59.4\%$ $70.0\%$ $61.7\%$ $54.3\%$ PertGraph $90.9\%$ $79.5\%$ $87.6\%$ $84.7\%$ $94.1\%$ $43.8\%$ $47.5\%$ $51.7\%$ $51.0\%$ $\epsilon=32$ $\mathsf{CARIBOU}$ $92.2\%$ $73.9\%$ $9.8\% 88$$8.2\%$ $95.8\%$ $90.6\%$ $82.5\%$ $78.3\%$ $67.0\%$ DPDGC $91.3\%$ $2.9\% 88$$8.8\%$ $79.9\%$ $94.3\%$ $43.8\%$ $50.0\%$ $51.7\%$ $40.0\%$ GAP $90.2\%$ $76.5\%$ $88.7\%$ $82.2\%$ $94.6\%$ $59.4\%$ $65.0\%$ $61.7\%$ $54.3\%$ PertGraph $90.6\%$ $79.9\%$ $86.9\%$ $85.2\%$ $94.4\%$ $43.8\%$ $47.5\%$ $51.7\%$ $51.0\%$ NDP (max node degree =20=20) $\epsilon=1$ $\mathsf{CARIBOU}$ $83.10\%$ $50.32\%$ $62.05\%$ $67.17\%$ $87.75\%$ $53.95\%$ $53.67\%$ $51.80\%$ $51.68\%$ DPDGC $54.42\%$ $36.73\%$ $51.34\%$ $30.04\%$ $42.60\%$ $48.06\%$ $47.25\%$ $52.20\%$ $48.99\%$ GAP $36.71\%$ $35.07\%$ $55.06\%$ $33.95\%$ $30.88\%$ $65.62\%$ $55.00\%$ $58.33\%$ $59.00\%$ PertGraph $23.91\%$ $18.58\%$ $37.32\%$ $13.62\%$ $17.10\%$ $51.88\%$ $51.25\%$ $49.83\%$ $50.60\%$ $\epsilon=2$ $\mathsf{CARIBOU}$ $84.60\%$ $53.03\%$ $64.22\%$ $69.63\%$ $88.89\%$ $54.49\%$ $54.32\%$ $52.23\%$ $52.14\%$ DPDGC $64.05\%$ $42.68\%$ $65.20\%$ $30.92\%$ $53.76\%$ $48.06\%$ $47.25\%$ $52.20\%$ $48.99\%$ GAP $42.32\%$ $37.52\%$ $61.44\%$ $3.67\% 33$$3.46\%$ $65.62\%$ $55.00\%$ $58.33\%$ $59.00\%$ PertGraph $31.32\%$ $20.39\%$ $38.08\%$ $14.6\%$ $19.90\%$ $51.88\%$ $51.25\%$ $49.83\%$ $50.60\%$ $\epsilon=4$ $\mathsf{CARIBOU}$ $85.95\%$ $55.73\%$ $67.15\%$ $71.97\%$ $89.89\%$ $4.71\% 55$$5.35\%$ $52.83\%$ $52.89\%$ DPDGC $70.79\%$ $47.52\%$ $78.82\%$ $31.22\%$ $71.17\%$ $48.06\%$ $47.25\%$ $52.20\%$ $48.99\%$ GAP $48.82\%$ $40.69\%$ $67.40\%$ $33.58\%$ $37.42\%$ $65.62\%$ $55.00\%$ $58.33\%$ $59.00\%$ PertGraph $35.14\%$ $22.50\%$ $39.03\%$ $16.40\%$ $23.48\%$ $51.88\%$ $51.25\%$ $49.83\%$ $50.60\%$ $\epsilon=8$ $\mathsf{CARIBOU}$ $87.15\%$ $58.05\%$ $70.17\%$ $73.80\%$ $90.73\%$ $55.71\%$ $56.09\%$ $53.57\%$ $53.50\%$ DPDGC $75.32\%$ $49.34\%$ $82.65\%$ $37.47\%$ $78.54\%$ $48.06\%$ $47.25\%$ $52.20\%$ $49.49\%$ GAP $53.75\%$ $42.60\%$ $71.14\%$ $33.07\%$ $45.18\%$ $65.62\%$ $55.00\%$ $58.33\%$ $59.00\%$ PertGraph $36.82\%$ $24.05\%$ $40.90\%$ $19.33\%$ $27.18\%$ $51.88\%$ $51.25\%$ $49.83\%$ $50.60\%$ $\epsilon=16$ $\mathsf{CARIBOU}$ $88.28\%$ $60.06\%$ $72.76\%$ $75.24\%$ $91.53\%$ $55.99\%$ $56.99\%$ $54.11\%$ $54.28\%$ DPDGC $77.85\%$ $50.00\%$ $84.27\%$ $53.71\%$ $81.84\%$ $48.06\%$ $47.25\%$ $52.20\%$ $50.00\%$ GAP $57.79\%$ $44.07\%$ $73.65\%$ $33.95\%$ $51.49\%$ $65.62\%$ $55.00\%$ $58.33\%$ $59.00\%$ PertGraph $39.57\%$ $25.95\%$ $43.45\%$ $22.14\%$ $31.64\%$ $51.88\%$ $51.25\%$ $49.83\%$ $50.60\%$ $\epsilon=32$ $\mathsf{CARIBOU}$ $89.35\%$ $61.87\%$ $74.83\%$ $76.61\%$ $92.36\%$ $56.23\%$ $57.72\%$ $54.64\%$ $54.77\%$ DPDGC $80.53\%$ $50.65\%$ $85.34\%$ $61.18\%$ $85.53\%$ $48.06\%$ $47.25\%$ $52.20\%$ $49.59\%$ GAP $60.98\%$ $45.17\%$ $75.54\%$ $37.79\%$ $56.28\%$ $65.62\%$ $55.00\%$ $58.33\%$ $59.00\%$ PertGraph $45.63\%$ $29.30\%$ $46.38\%$ $25.44\%$ $39.40\%$ $51.88\%$ $51.25\%$ $49.83\%$ $50.60\%$ Non-Private Plain $\mathsf{CARIBOU}$ $92.2\%$ $78.1\%$ $9.9\% 88$$8.4\%$ $95.8\%$ $\% 100.0100.0\%$ $\% 100.0100.0\%$ DPDGC $92.8\%$ $6.4\% 88$$8.1\%$ $83.9\%$ $96.2\%$ $59.4\%$ $77.5\%$ $63.3\%$ $73.0\%$ GAP $90.9\%$ $78.3\%$ $89.1\%$ $85.1\%$ $95.2\%$ $97.9\%$ $100.0\%$ $93.3\%$ $100.0\%$ PertGraph $91.6\%$ $79.7\%$ $87.0\%$ $82.1\%$ $94.2\%$ $9.4\% 55$$5.0\%$ $60.0\%$ $58.0\%$ MLP $83.9\%$ $51.1\%$ $84.9\%$ $72.9\%$ $91.1\%$ $40.6\%$ $47.5\%$ $46.7\%$ $51.0\%$
Figure 11: Classification Accuracy under $C_{\mathsf{L}},\alpha_{1},\beta$ of CGL.
A-G Larger Datasets Evaluation
To further assess the scalability on large real-world graphs, we additionally evaluate $\mathsf{CARIBOU}$ on the Reddit2 dataset [zeng2019graphsaint]. Reddit2 contains over 232K nodes and 23M edges, representing a significantly larger and structurally richer benchmark. Table XIV summarizes the top-1 accuracy under both EDP and NDP across privacy budgets $\epsilon\in\{1,2,4,8,16,32\}$.
TABLE XIV: Evaluating EDP and NDP over Reddit2 Dataset
$\mathbf{\epsilon}$ 1 2 4 8 16 32 EDP $70.6\%$ $73.8\%$ $76.4\%$ $78.6\%$ $78.7\%$ $80.9\%$ NDP $63.7\%$ $68.5\%$ $3.2\% 77$$7.3\%$ $78.4\%$ $79.6\%$
We observe: 1) $\mathsf{CARIBOU}$ achieves $70.6\%$ (EDP) and $63.7\%$ (NDP) accuracy even at $\epsilon=1$, improving steadily as privacy budgets relax, reaching $80.9\%$ (EDP) and $79.6\%$ (NDP) at $\epsilon=32$. This demonstrates that $\mathsf{CARIBOU}$ maintains stable utility even on million-edge graphs. 2) Due to the dataset’s large maximum degree and degree heterogeneity, NDP naturally incurs higher sensitivity. This is reflected in a larger EDP–NDP gap at small $\epsilon$, which narrows as noise reduces with increasing privacy budgets. 3) Despite the challenging properties of Reddit2, $\mathsf{CARIBOU}$ avoids the noise explosion potentially observed in non-contractive private GNNs. The monotonic accuracy increase across $\epsilon$ validates that $\mathsf{CARIBOU}$’s contractive operator bounds perturbation amplification over multiple hops. These results collectively demonstrate that CARIBOU extends effectively to large-scale real-world social networks, providing additional empirical support for the generality of $\mathsf{CARIBOU}$’s convergent privacy framework.
A-H Failure Case on Divergent Privacy
Table XV reports the accuracy of GAP under a fixed privacy budget $\epsilon=4$ as the depth $K$ increases. Since GAP relies on standard linear RDP composition, the required noise variance grows unboundedly with $K$, leading to a divergent noise allocation across layers. This phenomenon is clearly reflected in the performance trend: while GAP achieves $77.10\%$ accuracy at $K=2$, the accuracy steadily degrades as the GNN becomes deeper, dropping to $71.03\%$ at $K=64$. Figure 9 also confirms that $\mathsf{CARIBOU}$ shows better model accuracy than GAPacross different $K$. Moreover, GAPfurther drops to $51.11\%$ at $K=128$, approaching random-guessing performance. This study confirms that non-convergent privacy accounting inevitably leads to excessive noise at large depths, which severely impairs model utility.
TABLE XV: Accuracy with Divergent Noise Allocation. We set $\epsilon=4$ and instantiate GAP on the Coradataset as an example.
$K$ 22 44 88 $16$ $32$ $64$ $128$
Accuracy $77.10\%$ $75.65\%$ $74.91\%$ $73.43\%$ $74.17\%$ $71.03\%$ $51.11\%$
A-I More Related Works
A-I1 Differential privacy for graph structures